OpenStudy (anonymous):

What is the slope of the line passing through the points (2, 7) and (–1, 3)? 2/7 3/4 4/3 1/3

OpenStudy (johnweldon1993):

slope = (y2-y1)/(x2-x1) you have 2 points which is all you need (2, 7) and (–1, 3) (x1,y1) (x2,y2) plug those into that equation

OpenStudy (abb0t):

HINT: \(slope = m =\frac{ y_2-y_1 }{ x_2-x_1 }\)

OpenStudy (anonymous):

Thanks! i'll try to figure it out!

OpenStudy (anonymous):

cant find the right answer

OpenStudy (anonymous):

\[\frac{7-3}{2-(-1)}\]

OpenStudy (anonymous):

oooh..I see what i did wrong.